How many liters of a 50% alcohol solution must be mixed with 80 liters if a 90% solution to get a 70% solution ?
Answer by stanbon(75887)

How many liters of a 50% alcohol solution must be mixed with 80 liters if a 90% solution to get a 70% solution ?
------
Equation:
alc. + alc. = alc
0.50x + 0.90*80 = 0.70(x+80)
-----
50x + 90*80 = 70x + 70*80
20x = 20*80
x = 80 liters (amt. of 50% solution to add)
=======================
Cheers,
Stan H.
=======================